# One Homography is All You Need: IMM-based Joint Homography and Multiple Object State Estimation

### v3

@@ 2.4 Tracking Involving Homography
Various methods have used homography between multiple camera views to increase the robustness of the tracking of sports players. Since a particular player may be occluded in one camera view but not in another, using multiple cameras increases the robustness of a player tracking system to player occlusions. Examples of methods which make use of inter-camera homographies may be found in [65, 66, 67, 68, 69, 70]. However, few approaches consider the homography between a single camera view and the playing field. Notably, Hayet et al. [71] incrementally update the homography between the camera view and an imaged soccer field and explicitly use it to track player locations on the playing field with a Kalman filter. Recently, Maglo et al. [72] compute the homography to display player positions, but the tracking is performed without the aid of the homography. Instead, they rely heavily on appearance features extracted by a model fine-tuned offline after tracklet generation (on the corresponding bounding boxes). As a result, their method is not suitable for online operation. Most recently, Yi et al. [32] manually annotated the homographic projections between the image and ground planes of various datasets. They then perform tracking and association in the ground plane. However, as will be shown in Section 4, camera motion explicitly influences target position in their method – which this paper regards as a modelling error.
@@ 3 Background: Camera Motion Compensation and Motion Models
The majority of modern approaches to MOT model target motion in the image plane [8, 9, 10, 11, 12]. Particularly, the near-constant velocity model is usually employed:
@@
$\mathbf{x}^{I}_{0}$$\mathbf{y}^{I}_{0}$cdots $\mathbf{x}^{I}_{t}$$\mathbf{y}^{I}_{t}$
@@
and assumes that the camera remains stationary. As a result, several approaches [56, 52, 51, 55, 9, 10, 58, 61] make use of camera motion compensation after prediction with (1). Specifically, a matrix
@@
where the prime superscript denotes the camera-motion-compensated state component. $\mathbf{\hat{A}}_{t}$ is estimated with the camera calibration module of the OpenCV [73] library as in [10]. Most previous approaches make use of optical flow features and the RANSAC [74] algorithm, although [32] makes use of the enhanced correlation coefficient (ECC) optimisation method.
UCMCTrack [32] instead models target motion in the ground plane:
@@
where $\mathbf{x}^{W}_{t}=\begin{bmatrix}x^{W}_{t}&y^{W}_{t}\end{bmatrix}^{\top}$ represents the target’s x- and y-coordinates, respectively. The superscript $W$ indicates that the state is represented in the ground (or world) plane. The additive noise term $\mathbf{w}^{W}$ is sampled from a zero-mean multivariate normal distribution with covariance $\mathbf{Q}^{W,\mathbf{x}}$ (i.e. the process noise covariance matrix) determined by[32]:
@@
$\begin{bmatrix}x^{I}_{t}\\ x^{I}_{t}+0.5\times h^{I}_{t}\\ 1\end{bmatrix}=\operatorname{norm}\left(\mathbf{H}^{L}\begin{bmatrix}x^{W}_{t}\\ y^{W}_{t}\\ 1\end{bmatrix}\right),$ (8)
where $\operatorname{norm}(\begin{bmatrix}x&y&z\end{bmatrix}^{\top})=\begin{bmatrix}x/z&y/z&1\end{bmatrix}^{\top}$. For details regarding the theory of the pinhole camera model employed here, the reader is directed to [13, 75]. The left-hand side of (8) will subsequently be referred to as the bottom-centre bounding box coordinates. As in [32], these coordinates are selected to represent the projection of an object’s ground plane position since it is where the object is expected to touch the ground (e.g. if the object is a person, this coordinate represents where their feet touch the ground).
@@ 4 The Proposed Solution
Let $\mathbf{H}^{W}_{t}$ represent the projection matrix as it evolves over time. Its evolution may be described by [13]
@@
with $\mathbf{H}^{W}_{0}=\mathbf{H}^{L}$, and $\mathbf{\tilde{Q}}_{t}^{W,\mathbf{H}}$ is an additive noise term which is estimated as described in Section 4.3. The superscript $\mathbf{H}$ differentiates this noise term from that of the ground plane position components, and the tilde differentiates it from the static motion model introduced in (11). Equation 10 describes how the homographic projection changes as a function of the previous projection and the estimated camera motion. Importantly, this is independent of and does not influence target motion. Note that the derivation of (10) in [13] assumes that ground plane points are static. While this is not the case where the target is a dynamic object, the motion model is applied nonetheless, especially since the process noise estimation procedure in Section 4.3 dynamically adjusts the expected error statistics.
A graphical model which combines Equation 5 and Equation 10 with the measurement model in (8) is shown in Figure 3. Here, $\mathbf{x}^{I}_{t}$ represents the bottom-centre bounding box position of the target detection. According to Bayesian network theory, the target and camera projection dynamics are independent while the target bounding box is unobserved, but they become dependent when conditioned on $\mathbf{x}_{t}^{I}$ [76]. Thus, the temporal evolution of the ground plane and homography states are independent but are related through a common measurement in the image plane. This model may allow for a more accurate estimation of a target’s ground plane motion since its ground plane position prediction is not directly influenced by camera motion — only the position of its projection in the image plane.
$\mathbf{x}^{W}_{0}$$\mathbf{H}^{W}_{0}$$\mathbf{x}^{I}_{0}$cdots $\cdots$ $\mathbf{x}^{W}_{t}$$\mathbf{H}^{W}_{t}$$\mathbf{x}^{I}_{t}$
@@
The authors of [32] note that in scenes with relatively low camera motion, their method performs better without the camera motion compensation step in (9). This may be because the estimated affine matrix (2) is inaccurate, and noise is not considered in (9). To remedy this, the method proposed in this paper includes an additive noise term in (10) (see Section 4.1 and Section 4.3). Additionally, to better model the evolution of $\mathbf{H}^{W}_{t}$ when there is no camera motion, the following motion model is included:
@@
Furthermore, it has been shown that using buffered intersection over union (BIoU) during association can increase performance in scenes where the camera or the tracks are highly dynamic [12]. Besides, there may be times when the tracks leave the ground plane (for example, jumps or flips in the DanceTrack dataset [33]), causing their measured ground plane positions to fluctuate wildly. During these times, it may be better to perform association with an image-based association measure such as the BIoU. Even if tracks do not leave the ground plane, they may be occluded, which could cause partial detections that manifest similar behaviour to when tracks leave the ground plane. Nevertheless, considering bounding box dimensions during association increases performance, as shown in Section 6. Therefore, the proposed method uses the simple bounding box motion model proposed in [12]. To perform association, the ground plane and image-based association scores are taken into account proportionally according to the predicted model likelihood [77]:
@@
where $\mu_{t-1}^{j}$ denotes the probability that the target motion matches model $m^{j}$ at time $t-1$, $p_{ji}$ denotes the probability of a transition from model $m^{j}$ to model $m^{i}$, $r$ denotes the number of models under consideration and $\mathbf{x}^{I}_{1:t-1}$ represents all past observations. This is explained further in Section 4.4.2.
@@ 4.1 State Definition and Motion Models
The proposed method uses the constant velocity model for the ground plane dimensions as in (5). For the homography state elements, both the dynamics models of (10) and (11) are combined with an IMM filter. The reader is referred to [77, 78, 79, 80] for details regarding IMM filters, but an overview is provided in A. The process noise covariance matrix corresponding to the homography state elements is determined as Section 4.3 explains.
When a track’s position is predicted without a measurement update, it is said to be coasting, and it is in the coasted state. Otherwise, it is confirmed. In the image plane, the bounding box prediction is independent of the state vector in (13) while the track is confirmed. This paper makes use of the simple motion model presented in [12], which averages bounding box velocities over a buffer of previous measurements and adds this average to the previous bounding box measurement:
@@
$n$ denotes the size of the buffer, which stores the bounding box measurements $\mathbf{x}^{M}_{t-n:t-1}$ previously associated with the particular track. The superscript $M$ indicates a measurement, except in $\mathbf{x}^{M}_{t\mid t-1}$ which refers to the predicted measurement at time $t$ given the measurements from $t-n$ to $t-1$. Each measurement is stored as $\begin{bmatrix}l^{I}&t^{I}&r^{I}&b^{I}\end{bmatrix}$, where the components denote the left x-, top y-, right x- and bottom y-coordinate of the corresponding bounding box, respectively. The maximum size of the bounding box measurement buffer is fixed to $n=5$. When fewer than two measurements are in the buffer, the last-known bounding box is used as the prediction.
@@ 4.2 Measurement Model
[…]
@@
$\begin{split}&\frac{\partial\left(x^{I}_{t},y^{I}_{t}\right)}{\partial\left(h_{1},h_{4},h_{7},h_{2},h_{5},h_{8},h_{3},h_{6}\right)}=\\ &\gamma\begin{bmatrix}x^{W}_{t}&0&-x^{I}_{t}x^{W}_{t}&y^{W}_{t}&0&-x^{I}_{t}y^{W}_{t}&1&0\\ 0&x^{W}_{t}&-y^{I}_{t}x^{W}_{t}&0&y^{W}_{t}&-y^{I}_{t}y^{W}_{t}&0&1\end{bmatrix}.\end{split}$ (20)
In practice, $\mathbf{H}^{W}_{t}$ is normalised with respect to $h_{9}$ such that $h_{9}=1$. Therefore, the partial derivative with respect to $h_{9}$ is 00.
@@ 4.3 Dynamic Noise Estimation
Instead of attempting to determine a static process noise covariance matrix for the dynamics models in (10) and (11), the proposed method obtains dynamic estimates of these parameters. The same is performed for the measurement noise covariance matrix associated with the measurement model (15) – this is shown to increase performance in Table 1.
From [81], consider a generalised representation of the process noise covariance matrix $\mathbf{Q}_{t}$, which can represent either the static motion model of (11), or the dynamic motion model of (10). This matrix is the expected value of the outer product of the state process noise vector $\mathbf{w}_{t}$:
@@
When a measurement update is performed, the measurement noise covariance matrix for the current time step is estimated by [81]:
@@
After $\mathbf{R}_{t}$ is obtained from (24), the resulting Kalman gain $\mathbf{K}_{t}$ is used to obtain an estimate of the process noise covariance matrix for the current time step [81]:
@@ 4.4 Association and Track Management
This section describes the association algorithm and how the BIoU and Mahalanobis distances are combined to form robust association scores between tracks and candidate detections. The Hungarian algorithm (specifically, the Jonker-Volgenant variation [82]) uses these scores in cascaded matching stages to obtain track-detection associations. There are three association stages, with association score thresholds $\alpha_{1},\alpha_{2}$ and $\alpha_{3}$, respectively. The values of the hyperparameters introduced in this section are obtained as described in Section 5.
@@ 4.4.2 The Association Algorithm
Figure 6 depicts the association algorithm graphically. At the start of a time step, the detections for the corresponding video frame are obtained, and all existing tracks’ positions are predicted by their ground- and image-plane filters. From here, confirmed and coasted tracks are considered by two cascaded association stages. Before the first stage, detections are separated into high- and low-confidence detections by the confidence thresholds $d_{\text{high}}$ and $d_{\text{low}}$, as introduced in [9]. Specifically, detections with confidence greater than or equal to $d_{\text{high}}$ are considered high-confidence, and those with confidence less than $d_{\text{high}}$ but greater than or equal to $d_{\text{low}}$, are considered low-confidence. Detections with confidences below $d_{\text{low}}$ are discarded.
@@
where $d_{\text{conf}}$ refers to the detection confidence value output by the detector, and $D$ is the normalised Mahalanobis distance as in [32]:
@@
where $\operatorname{CDF}$ is the cumulative distribution function of the chi-squared distribution, which is defined to have $24$ degrees of freedom (determined experimentally). As a result, $P(D)$ is in the range $[0,1]$. Finally, $\text{BIoU}(\mathbf{x}^{M}_{t},\mathbf{x}^{M}_{t|t-1})$ denotes the BIoU between a given detection $\mathbf{x}^{M}_{t}$ and the bounding box predicted by (14). The BIoU is calculated as in [12], where the IoU is calculated after the measured and predicted bounding box widths and heights are each scaled by $12b+$ 1; in this paper, only a single buffer scale parameter, $b$, is used.
@@
This association stage manifests the filter architecture depicted in Figure 5. While the ground plane filters are mixed explicitly with an IMM filter as established in the literature [80, 77, 78, 79] and explained in A, the image and ground plane filters are handled independently. Transition probabilities between these filters are only defined to use the predicted model likelihood (12) during association. Each model probability is initialised as $\mu^{I}_{0}=\mu^{W}_{0}=0.5$. When a measurement is associated with a track, these model probabilities are updated [77]:
@@
The second association stage helps to avoid tracks from coasting when their detections in the image plane diverge from their expected value as predicted by the ground plane filter. This may occur due to partial occlusion or irregular, rapid object motion. It may also occur when ground plane filter state estimates are inaccurate, notably when a track is re-identified after a period during which it was coasted [11]. In these cases, the predicted model likelihood $\mu^{W}_{t|t-1}$ is expected to be close to zero, while the predicted model likelihood $\mu^{I}_{t|t-1}$ is expected to be close to one. Thus, the association score (30) will be dominated by the BIoU, and the track may still be associated despite low $P(D)$. Note that [11] introduced a re-update step to mitigate the effect of state estimation noise and that the second association stage proposed here could be considered an alternative to their solution.
The image and ground plane filters are not mixed explicitly because a target’s ground plane state estimate should not be significantly influenced when the BIoU dominates the association score since this often implies motion away from the ground plane (e.g. a jump) or a partial detection, both of which are expected to result in an inaccurate measurement of the ground plane position. However, if the primary goal is to reduce the effect of state estimation noise as in [11], mixing ground and image plane state estimates may be beneficial.

### v4

@@ 2.4 Tracking Involving Homography
Various methods have used homography between multiple camera views to increase the robustness of the tracking of sports players. Since a particular player may be occluded in one camera view but not in another, using multiple cameras increases the robustness of a player tracking system to player occlusions. Examples of methods which make use of inter-camera homographies may be found in (Iwase2003TrackingViews; Iwase2004ParallelImages; Eshel2008HomographyCrowd; Seo2009HumanHomography; Sainan2018AthletesStudy; Khan2006AConstraint). However, few approaches consider the homography between a single camera view and the playing field. Notably, Hayet et al. (Hayet2005ATracking) incrementally update the homography between the camera view and an imaged soccer field and explicitly use it to track player locations on the playing field with a Kalman filter. Recently, Maglo et al. (Maglo2023IndividualView) compute the homography to display player positions, but the tracking is performed without the aid of the homography. Instead, they rely heavily on appearance features extracted by a model fine-tuned offline after tracklet generation (on the corresponding bounding boxes). As a result, their method is not suitable for online operation. Most recently, Yi et al. (ucmc) manually annotated the homographic projections between the image and ground planes of various datasets. They then perform tracking and association in the ground plane. However, as will be shown in Section 4, camera motion explicitly influences target position in their method – which this paper regards as a modelling error.
@@ 3 Background: Camera Motion Compensation and Motion Models
The majority of modern approaches to MOT model target motion in the image plane (Bewley2016SimpleTracking; Zhang2022ByteTrack:Box; botsort; Cao2022Observation-CentricTracking; Yang2022HardSpace). Particularly, the near-constant velocity model is usually employed:
@@
$\mathbf{x}^{I}_{0}$$\mathbf{y}^{I}_{0}$ $\cdots$ $\mathbf{x}^{I}_{t}$$\mathbf{y}^{I}_{t}$
@@
and assumes that the camera remains stationary. As a result, several approaches (bells; invisible; modelling_ambiguous; giao; Zhang2022ByteTrack:Box; botsort; Wang2022SMILEtrack:Tracking; interframe) make use of camera motion compensation after prediction with (1). Specifically, a matrix
@@
where the prime superscript denotes the camera-motion-compensated state component. $\mathbf{\hat{A}}_{t}$ is estimated with the camera calibration module of the OpenCV (opencv) library as in (botsort). Most previous approaches make use of optical flow features and the RANSAC (ransac) algorithm, although (ucmc) makes use of the enhanced correlation coefficient (ECC) optimisation method.
UCMCTrack (ucmc) instead models target motion in the ground plane:
@@
where $\mathbf{x}^{W}_{t}=\begin{bmatrix}x^{W}_{t}&y^{W}_{t}\end{bmatrix}^{\top}$ represents the target’s x- and y-coordinates, respectively. The superscript $W$ indicates that the state is represented in the ground (or world) plane. The additive noise term $\mathbf{w}^{W}$ is sampled from a zero-mean multivariate normal distribution with covariance $\mathbf{Q}^{W,\mathbf{x}}$ (i.e. the process noise covariance matrix) determined by(ucmc):
@@
[…]
where $\operatorname{norm}(\begin{bmatrix}x&y&z\end{bmatrix}^{\top})=\begin{bmatrix}x/z&y/z&1\end{bmatrix}^{\top}$. For details regarding the theory of the pinhole camera model employed here, the reader is directed to (bhitk; multiple_view_geometry). The left-hand side of (8) will subsequently be referred to as the bottom-centre bounding box coordinates. As in (ucmc), these coordinates are selected to represent the projection of an object’s ground plane position since it is where the object is expected to touch the ground (e.g. if the object is a person, this coordinate represents where their feet touch the ground).
@@ 4 The Proposed Solution
Let $\mathbf{H}^{W}_{t}$ represent the projection matrix as it evolves over time. Its evolution may be described by (bhitk)
@@
with $\mathbf{H}^{W}_{0}=\mathbf{H}^{L}$, and $\mathbf{\tilde{Q}}_{t}^{W,\mathbf{H}}$ is an additive noise term which is estimated as described in Section 4.3. The superscript $\mathbf{H}$ differentiates this noise term from that of the ground plane position components, and the tilde differentiates it from the static motion model introduced in (11). Equation 10 describes how the homographic projection changes as a function of the previous projection and the estimated camera motion. Importantly, this is independent of and does not influence target motion. Note that the derivation of (10) in (bhitk) assumes that ground plane points are static. While this is not the case where the target is a dynamic object, the motion model is applied nonetheless, especially since the process noise estimation procedure in Section 4.3 dynamically adjusts the expected error statistics.
A graphical model which combines Equation 5 and Equation 10 with the measurement model in (8) is shown in Figure 3. Here, $\mathbf{x}^{I}_{t}$ represents the bottom-centre bounding box position of the target detection. According to Bayesian network theory, the target and camera projection dynamics are independent while the target bounding box is unobserved, but they become dependent when conditioned on $\mathbf{x}_{t}^{I}$ (BIS). Thus, the temporal evolution of the ground plane and homography states are independent but are related through a common measurement in the image plane. This model may allow for a more accurate estimation of a target’s ground plane motion since its ground plane position prediction is not directly influenced by camera motion — only the position of its projection in the image plane.
$\mathbf{x}^{W}_{0}$$\mathbf{H}^{W}_{0}$$\mathbf{x}^{I}_{0}$ $\cdots$ $\cdots$ $\mathbf{x}^{W}_{t}$$\mathbf{H}^{W}_{t}$$\mathbf{x}^{I}_{t}$
@@
The authors of (ucmc) note that in scenes with relatively low camera motion, their method performs better without the camera motion compensation step in (9). This may be because the estimated affine matrix (2) is inaccurate, and noise is not considered in (9). To remedy this, the method proposed in this paper includes an additive noise term in (10) (see Section 4.1 and Section 4.3). Additionally, to better model the evolution of $\mathbf{H}^{W}_{t}$ when there is no camera motion, the following motion model is included:
@@
Furthermore, it has been shown that using buffered intersection over union (BIoU) during association can increase performance in scenes where the camera or the tracks are highly dynamic (Yang2022HardSpace). Besides, there may be times when the tracks leave the ground plane (for example, jumps or flips in the DanceTrack dataset (dance)), causing their measured ground plane positions to fluctuate wildly. During these times, it may be better to perform association with an image-based association measure such as the BIoU. Even if tracks do not leave the ground plane, they may be occluded, which could cause partial detections that manifest similar behaviour to when tracks leave the ground plane. Nevertheless, considering bounding box dimensions during association increases performance, as shown in Section 6. Therefore, the proposed method uses the simple bounding box motion model proposed in (Yang2022HardSpace). To perform association, the ground plane and image-based association scores are taken into account proportionally according to the predicted model likelihood (imm):
@@
where $\mu_{t-1}^{j}$ denotes the probability that the target motion matches model $m^{j}$ at time t1t-1, $p_{ji}$ denotes the probability of a transition from model $m^{j}$ to model $m^{i}$, $r$ denotes the number of models under consideration and $\mathbf{x}^{I}_{1:t-1}$ represents all past observations. This is explained further in Section 4.4.2.
@@ 4.1 State Definition and Motion Models
The proposed method uses the constant velocity model for the ground plane dimensions as in (5). For the homography state elements, both the dynamics models of (10) and (11) are combined with an IMM filter to remain robust in both static and dynamic camera motion scenarios. The reader is referred to (imm; Blackman1999DesignSystems; imm_genovese; Blom1988TheCoefficients) for details regarding IMM filters, but an overview is provided in A. The process noise covariance matrix corresponding to the homography state elements is determined as Section 4.3 explains.
When a track’s position is predicted without a measurement update, it is said to be coasting, and it is in the coasted state. Otherwise, it is confirmed. In the image plane, the bounding box prediction is independent of the state vector in (13) while the track is confirmed. This paper makes use of the simple motion model presented in (Yang2022HardSpace), which averages bounding box velocities over a buffer of previous measurements and adds this average to the previous bounding box measurement:
@@
$n$ denotes the size of the buffer, which stores the bounding box measurements $\mathbf{x}^{M}_{t-n:t-1}$ previously associated with the particular track. The superscript $M$ indicates a measurement, except in $\mathbf{x}^{M}_{t\mid t-1}$ which refers to the predicted measurement at time $t$ given the measurements from tnt-n to t1t-1. Each measurement is stored as $\begin{bmatrix}l^{I}&t^{I}&r^{I}&b^{I}\end{bmatrix}$, where the components denote the left x-, top y-, right x- and bottom y-coordinate of the corresponding bounding box, respectively. The maximum size of the bounding box measurement buffer is fixed to $n=5$. When fewer than two measurements are in the buffer, the last-known bounding box is used as the prediction.
@@ 4.2 Measurement Model
$\mathbf{b}=\begin{bmatrix}b_{1}\\ b_{2}\\ b_{3}\end{bmatrix}=\mathbf{H}^{W}_{t}\begin{bmatrix}x^{W}_{t}\\ y^{W}_{t}\\ 1\end{bmatrix}$ (17)
@@
[…]
In practice, $\mathbf{H}^{W}_{t}$ is normalised with respect to $h_{9}$ such that $h_{9}=1$. Therefore, the partial derivative with respect to $h_{9}$ is 0.
@@ 4.3 Dynamic Noise Estimation
Instead of attempting to determine a static process noise covariance matrix for the dynamics models in (10) and (11), the proposed method obtains dynamic estimates of these parameters. This allows IMM-JHSE to exploit the possibility of different targets having different noise parameters, and allows these estimates to be robust against irregularities such as when targets leave the ground plane or occlude one another (often resulting in partial detections). The same is performed for the measurement noise covariance matrix associated with the measurement model (15) – this is shown to increase performance in Table 1.
From (adaptive_noise), consider a generalised representation of the process noise covariance matrix $\mathbf{Q}_{t}$, which can represent either the static motion model of (11), or the dynamic motion model of (10). This matrix is the expected value of the outer product of the state process noise vector $\mathbf{w}_{t}$:
@@
When a measurement update is performed, the measurement noise covariance matrix for the current time step is estimated by (adaptive_noise):
@@
After $\mathbf{R}_{t}$ is obtained from (24), the resulting Kalman gain $\mathbf{K}_{t}$ is used to obtain an estimate of the process noise covariance matrix for the current time step (adaptive_noise):
@@
4.3.1 Convergence
Dynamic noise estimates are obtained independently for each track; it is therefore not sufficient to show empirical evidence of convergence for a small number of select tracks to prove convergence in general, yet an investigation of convergence over all tracks is complicated by the possibility that different targets have different true noise covariance matrices. Furthermore, because target motion may be intermittently irregular, the true noise covariance matrices are not expected to be static. We refer to (adaptive_noise) for the original derivation of the above formulae and for their analysis of filter convergence. Furthermore, under the interpretation that the noise estimates are obtained by a sample covariance estimator, the Cramér-Rao lower bound can be shown to be monotonically decreasing as the window size increases (Kay1993). Therefore, convergence is theoretically possible with the assumption that the true noise covariance is static.
The rigorous analysis of the convergence of these estimates is deferred to future work.
@@ 4.4 Association and Track Management
This section describes the association algorithm and how the BIoU and Mahalanobis distances are combined to form robust association scores between tracks and candidate detections. The Hungarian algorithm (specifically, the Jonker-Volgenant variation (lapjv)) uses these scores in cascaded matching stages to obtain track-detection associations. There are three association stages, with association score thresholds $\alpha_{1},\alpha_{2}$ and $\alpha_{3}$, respectively. The values of the hyperparameters introduced in this section are obtained as described in Section 5.
@@ 4.4.2 The Association Algorithm
Figure 6 depicts the association algorithm graphically. At the start of a time step, the detections for the corresponding video frame are obtained, and all existing tracks’ positions are predicted by their ground- and image-plane filters. From here, confirmed and coasted tracks are considered by two cascaded association stages. Before the first stage, detections are separated into high- and low-confidence detections by the confidence thresholds $d_{\text{high}}$ and $d_{\text{low}}$, as introduced in (Zhang2022ByteTrack:Box). Specifically, detections with confidence greater than or equal to $d_{\text{high}}$ are considered high-confidence, and those with confidence less than $d_{\text{high}}$ but greater than or equal to $d_{\text{low}}$, are considered low-confidence. Detections with confidences below $d_{\text{low}}$ are discarded.
@@
where $d_{\text{conf}}$ refers to the detection confidence value output by the detector, and $D$ is the normalised Mahalanobis distance as in (ucmc):
@@
where $\operatorname{CDF}$ is the cumulative distribution function of the chi-squared distribution, which is defined to have $24$ degrees of freedom (determined experimentally). As a result, $P(D)$ is in the range $[0,1]$. Finally, $\text{BIoU}(\mathbf{x}^{M}_{t},\mathbf{x}^{M}_{t|t-1})$ denotes the BIoU between a given detection $\mathbf{x}^{M}_{t}$ and the bounding box predicted by (14). The BIoU is calculated as in (Yang2022HardSpace), where the IoU is calculated after the measured and predicted bounding box widths and heights are each scaled by $2b+1$; in this paper, only a single buffer scale parameter, $b$, is used.
@@
This association stage manifests the filter architecture depicted in Figure 5. While the ground plane filters are mixed explicitly with an IMM filter as established in the literature (Blom1988TheCoefficients; imm; Blackman1999DesignSystems; imm_genovese) and explained in A, the image and ground plane filters are handled independently. Transition probabilities between these filters are only defined to use the predicted model likelihood (12) during association. Each model probability is initialised as $\mu^{I}_{0}=\mu^{W}_{0}=0.5$. When a measurement is associated with a track, these model probabilities are updated (imm):
@@
The second association stage helps to avoid tracks from coasting when their detections in the image plane diverge from their expected value as predicted by the ground plane filter. This may occur due to partial occlusion or irregular, rapid object motion. It may also occur when ground plane filter state estimates are inaccurate, notably when a track is re-identified after a period during which it was coasted (Cao2022Observation-CentricTracking). In these cases, the predicted model likelihood $\mu^{W}_{t|t-1}$ is expected to be close to zero, while the predicted model likelihood $\mu^{I}_{t|t-1}$ is expected to be close to one. Thus, the association score (30) will be dominated by the BIoU, and the track may still be associated despite low $P(D)$. Note that (Cao2022Observation-CentricTracking) introduced a re-update step to mitigate the effect of state estimation noise and that the second association stage proposed here could be considered an alternative to their solution.
The image and ground plane filters are not mixed explicitly because a target’s ground plane state estimate should not be significantly influenced when the BIoU dominates the association score since this often implies motion away from the ground plane (e.g. a jump) or a partial detection, both of which are expected to result in an inaccurate measurement of the ground plane position. However, if the primary goal is to reduce the effect of state estimation noise as in (Cao2022Observation-CentricTracking), mixing ground and image plane state estimates may be beneficial.
@@
The hybrid association strategy outlined above is expected to afford IMM-JHSE enhanced robustness to situations where targets unexpectedly leave the ground plane or undergo irregular motion that would otherwise cause their motion to diverge from the prescribed ground-plane motion model.